From the information given here, determine the 95% confidence interval estimate of the population mean.
x(bar) = 20 σ= 18 n = 36
Answer #1

Confidence interval for Population mean is given as below:

Confidence interval = Xbar ± Z*σ/sqrt(n)

From given data, we have

Xbar = 20

σ = 18

n = 36

Confidence level = 95%

Critical Z value = 1.96

(by using z-table)

Confidence interval = Xbar ± Z*σ/sqrt(n)

Confidence interval = 20 ± 1.96*18/sqrt(36)

Confidence interval = 20 ± 5.8799

Lower limit = 20 - 5.8799 = 14.12

Upper limit = 20 + 5.8799 = 25.88

Confidence interval = (14.12, 25.88)